# CtRL-Sim: Reactive and Controllable Driving Agents with Offline Reinforcement Learning

### v1

@@ Appendix I CtRL-Sim Finetuning
The Waymo Open Motion dataset largely contains nominal driving scenes. To enhance control over the generation of safety-critical scenarios, we finetune CtRL-Sim on a simulated dataset of safety-critical scenarios generated by CtRL-Sim. We first collect a simulated dataset of scenes either containing a vehicle-vehicle collision or an offroad infraction. Specifically, we generate rollouts of a single agent with the negatively tilted base CtRL-Sim model where the other agents are set to log replay through physics, and we save the scenario only if the generated rollout yields a vehicle-vehicle collision or vehicle-road-edge collision. For tilting, we uniformly sample $\kappa_{\text{veh-veh}}\sim\mathcal{U}(-25,0)$ and $\kappa_{\text{goal}}\sim\mathcal{U}(-25,0)$ when generating vehicle-vehicle collision scenarios, and we uniformly sample $\kappa_{\text{veh-edge}}\sim\mathcal{U}(-25,0)$ and $\kappa_{\text{goal}}\sim\mathcal{U}(-25,0)$ when generating vehicle-road-edge collision scenarios. By additionally negatively tilting the goal, this grants the model more flexibility when generating traffic violations as the agents are not trying to reach its prescribed goal. We collect 5000 scenarios of each type of traffic violation derived from the training set, which comprises the simulated dataset of safety-critical scenarios. To encourage CtRL-Sim to learn how to generate safety-critical scenarios without forgetting how to generate good driving behaviour, we adopt a continual pre-training strategy for finetuning (Ibrahim et al., 2024) where we randomly sample 90000 real training scenarios from the offline RL dataset in each training epoch, or a 90% replay ratio. We rewarm the learning rate to the maximum learning rate of 5e-4 over 500 steps and follow a linear decay learning rate schedule to 0 over 20 epochs. We expect that the finetuned CtRL-Sim model will be more capable of generating long-tail scenarios as it is more exposed to such scenarios during finetuning.
@@
Figure 7 shows more qualitative examples demonstrating the effects of positive exponential tilting on each of the three reward components. In the left panels, CtRL-Sim with no tilting produces a vehicle-vehicle collision between two interacting agents at a left-turn. With positive vehicle-vehicle tilting, the CtRL-Sim-controlled agent moves more to the right-hand side of the lane to avoid the collision. In the middle panels, CtRL-Sim with no tilting produces a vehicle-edge collision as the bus pulls into the curb. With positive vehicle-edge tilting, the CtRL-Sim-controlled agent pulls into the curb at a safer distance from the curb. In the right panels, CtRL-Sim with no tilting reaches the goal. With positive goal tilting, the CtRL-Sim-controlled agent reaches the goal much faster and nearly avoids collision with the turning vehicle. In Figure 8, we show two more examples of adversarial collision scenarios generated with negative vehicle-vehicle tilting. We refer the interested reader to the supplementary materials for more example videos.

### v2

@@ Appendix H CAT Simulated Data Collection and CtRL-Sim Finetuning
To encourage CtRL-Sim to learn how to generate safety-critical scenarios without forgetting how to generate good driving behaviour, we adopt a continual pre-training strategy for finetuning [46] where we randomly sample 3577 real training scenarios from the offline RL dataset in each training epoch, or a 50% replay ratio. We rewarm the learning rate to the maximum learning rate of 5e-4 over 500 steps and follow a linear decay learning rate schedule to 0 over 20 epochs. We expect that the finetuned CtRL-Sim model will be more capable of generating long-tail scenarios as it is more exposed to such scenarios during finetuning. Finetuning takes roughly 30 minutes on 1 NVIDIA A100-Large GPU.
Appendix I Adversarial Scenario Generation User Study
We first conduct a pilot study with 5 participants to evaluate which method (CtRL-Sim vs. CAT) generates more plausible adversarial behaviours. Each participant was tasked with evaluating the same set of 25 pairs of scenarios that were randomly selected from a pool of 259 scenarios, where one scenario employed a CtRL-Sim adversarial agent and the other was a CAT adversarial agent. In both instances, the planner was a positively tilted CtRL-Sim planner. Table 6 shows the tally of the votes, where we observed strong preference for adversarial scenarios generated using CtRL-Sim.
Method Times Preferred
CtRL-Sim 97
CAT 96
Tie 66
Table 7: The tally of votes for the larger study. We show the breakdown of the votes for the conducted larger study. We observe that the results of this study are less conclusive than the pilot study.
We conduct a larger study including a total of $22$ participants. We did not record any identifying information of the participants, and participants were invited on a voluntary basis. Users were presented with the user interface shown in Figure 6, where we randomize the order between both videos. Each participant responded to a total of $30$ pairs of scenarios, which took an average $12$ minutes to complete. These pairs were randomly selected from the same pool of $259$ scenarios as the pilot study. Table 7 shows the tally of this larger study. As we can see, the results of this larger study are inconclusive compared to the pilot study. We hypothesize that the framing of the question may have been ambiguous for a larger audience, and that the pairs of videos may have included less interesting scenarios (e.g, adversary rear-ending planner), resulting in many comparisons that were very difficult to judge, thus adding noise to the survey results. We plan to further study the question of plausibility in a future user study with better scenario selection.
Figure 6: User Study Example Scenario. We show an example of a pair of scenarios along with the question users are asked to answer.
@@ Appendix I Adversarial Scenario Generation User Study
Figure 7 shows more qualitative examples demonstrating the effects of positive exponential tilting on each of the three reward components. In the left panels, CtRL-Sim with no tilting produces a vehicle-vehicle collision between two interacting agents at a left-turn. With positive vehicle-vehicle tilting, the CtRL-Sim-controlled agent moves more to the right-hand side of the lane to avoid the collision. In the middle panels, CtRL-Sim with no tilting produces a vehicle-edge collision as the bus pulls into the curb. With positive vehicle-edge tilting, the CtRL-Sim-controlled agent pulls into the curb at a safer distance from the curb. In the right panels, CtRL-Sim with no tilting reaches the goal. With positive goal tilting, the CtRL-Sim-controlled agent reaches the goal much faster and nearly avoids collision with the turning vehicle. In Figure 8, we show two more examples of adversarial collision scenarios generated with negative vehicle-vehicle tilting. We refer the interested reader to the supplementary video for more examples.
@@
Appendix K Multi-Agent Simulation Results with Higher Temperature Sampling
In Table 8, we report results from the same experiments as Table 1 except with a higher action sampling temperature, set to $1.5$.
Reconstruction Distributional Realism Common Sense
[…]
Actions-Only [9] 11.70 $\pm$ 1.12 4.78 $\pm$ 0.42 34.4 $\pm$ 1.3 14.3 $\pm$ 0.3 22.8 $\pm$ 0.7 29.7 $\pm$ 1.7
Imitation Learning 2.42 $\pm$ 0.17 1.47 $\pm$ 0.07 73.8 $\pm$ 1.2 12.3 $\pm$ 0.5 7.3 $\pm$ 0.6 13.1 $\pm$ 0.4
DT (Max Return) [15] 3.25 $\pm$ 0.17 1.67 $\pm$ 0.05 60.5 $\pm$ 1.2 12.3 $\pm$ 0.4 6.1 $\pm$ 0.7 11.6 $\pm$ 0.3
CtRL-Sim (No State Prediction) 2.57 $\pm$ 0.16 1.52 $\pm$ 0.07 66.2 $\pm$ 1.0 12.3 $\pm$ 0.3 7.6 $\pm$ 0.7 13.1 $\pm$ 0.3
CtRL-Sim (Base) 2.49 $\pm$ 0.10 1.50 $\pm$ 0.04 67.9 $\pm$ 1.2 12.2 $\pm$ 0.2 7.6 $\pm$ 0.3 13.1 $\pm$ 0.5
CtRL-Sim (Positive Tilting) 2.38 $\pm$ 0.08 1.44 $\pm$ 0.03 67.2 $\pm$ 1.0 12.1 $\pm$ 0.1 6.7 $\pm$ 0.4 12.3 $\pm$ 0.3
$\text{DT}^{*}$ (GT Initial Return) 1.94 $\pm$ 0.07 1.28 $\pm$ 0.02 73.7 $\pm$ 1.5 12.2 $\pm$ 0.3 6.6 $\pm$ 0.4 12.6 $\pm$ 0.4
[…]
Table 8: Multi-agent simulation results over 1000 test scenes with action temperature = 1.5 over 3 seeds. This table presents the results from the same experiments as Table 1, but with an action sampling temperature of 1.5 instead of 1.0. This allows for a comparison of the impact of the temperature hyperparameter. Overall, an action sampling temperature of 1.0 yields better results.
Appendix L Fine-tuning CtRL-Sim on CtRL-Sim Scenarios
Figure 9: Effects of exponential tilting. Comparison of CtRL-Sim base model (magenta) and a CtRL-Sim model fine-tuned on adversarial CtRL-Sim scenarios (purple). As opposed to Figure 4, this fine-tuned model does not involve using CAT to select the adversarial scenarios. Rewards range from -25 to 25 for vehicle-vehicle collision (left), vehicle-edge collision (middle), and goal reaching (right). Results show smooth controllability, with fine-tuning enhancing this effect. Mean and std are reported over 5 seeds.
Instead of finetuning on CAT scenarios, we explore finetuning CtRL-Sim on adversarial scenarios generated by CtRL-Sim. We first collect a simulated dataset of scenes either containing a vehicle-vehicle collision or an offroad infraction. Specifically, we generate rollouts of a single agent with the negatively tilted base CtRL-Sim model where the other agents are set to log replay through physics, and we save the scenario only if the generated rollout yields a vehicle-vehicle collision or vehicle-road-edge collision. For tilting, we uniformly sample $\kappa_{\text{veh-veh}}\sim\mathcal{U}(-25,0)$ and $\kappa_{\text{goal}}\sim\mathcal{U}(-25,0)$ when generating vehicle-vehicle collision scenarios, and we uniformly sample $\kappa_{\text{veh-edge}}\sim\mathcal{U}(-25,0)$ and $\kappa_{\text{goal}}\sim\mathcal{U}(-25,0)$ when generating vehicle-road-edge collision scenarios. By additionally negatively tilting the goal, this grants the model more flexibility when generating traffic violations as the agents are not trying to reach its prescribed goal. We collect 5000 scenarios of each type of traffic violation derived from the training set, which comprises the simulated dataset of safety-critical scenarios. To encourage CtRL-Sim to learn how to generate safety-critical scenarios without forgetting how to generate good driving behaviour, we adopt the same finetuning strategy as in Appendix H, except we randomly sample 90000 real training scenarios from the offline RL dataset in each training epoch, or a 90% replay ratio. We find it useful to use a larger replay ratio when finetuning on CtRL-Sim scenarios. The controllability results are shown in Figure 9, demonstrating similar control over adversarial behaviours as the CAT-finetuned CtRL-Sim model.